# BlockPruner: Fine-grained Pruning for Large Language Models

### v3

@@ Appendix C Details of Evaluations
To ensure a fair and comprehensive comparison, we employed the same version of the LM Evaluation Harness as used in the SliceGPT experiments and obtained evaluation scores under identical experimental configurations. These scores closely match those reported in the SliceGPT paper, as detailed in Table 3. For consistency, we present our reproduced results in the main experiments.
@@
To evaluate the performance of pruned models on downstream tasks, we utilized five multiple-choice QA datasets: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c. Additionally, to assess text generation quality, we calculated perplexity using the test set of the Wikitext2 dataset. For the downstream task evaluations, we adhered to the default evaluation parameters and zero-shot settings, with a batch size set to 1. For perplexity calculations, the maximum text length was set to 2048, maintaining a batch size of 1 as well.
Appendix D Ablation Experiments under Higher Sparsity
The motivation behind BlockPruner is to maintain model performance more effectively through fine-grained block pruning. Evaluating the performance differences of block pruning at varying granularities under higher sparsity is essential, as demonstrating effectiveness under such conditions would further corroborate our motivations and claims. In light of this, we conducted ablation experiments with higher sparsity ratios on LLaMA-2-7B and LLaMA-2-13B models and found that our motivation and claims remain valid. The results are presented in Table 4.
@@ Appendix D Ablation Experiments under Higher Sparsity
Appendix E Inference Speed after Pruning
Since our method reduces more MHA blocks at the same pruning ratio compared to other methods, it consequently reduces the usage of the KV cache at equivalent pruning ratios. This reduction can potentially accelerate the inference speed of LLMs. To verify that our method indeed accelerates the model’s inference process, we compared the time taken to generate 128 tokens by models obtained from different pruning methods with KV caches. To ensure the robustness of the experimental results, we repeated each configuration 20 times and took the average inference time as the result. As observed from the Table 5, our method achieves the highest speedup at comparable pruning ratios.
@@ Appendix E Inference Speed after Pruning
Appendix F Time Costs of Various Pruning Methods
@@
Appendix G Post-training after Pruning
We sampled 8,000 instances from the Alpaca dataset and performed simple post-training on the pruned Llama2-7B and Llama2-13B obtained through BlockPruner using LoRA. We trained all linear layers except for the embedding layer and the language model head. The parameters LoRA rank and LoRA $\alpha$ were set to 32 and 10, respectively, with a learning rate of 2e-4 and a training batch size of 1. We also config the gradient accumulation steps to 4 and use a linear learning rate scheduler. We controlled the pruning ratios within the range of 24% to 33%. The results are shown in Figure 8. It can be seen that after training, our models showed further improvement at different pruning ratios. The Llama2-7B and Llama2-13B models recovered to 89% and 92% of the performance of the unpruned models, respectively, when pruned by approximately 1/4.
@@ Appendix G Post-training after Pruning
Appendix H Sensitivity to Sample Size
@@
Appendix I Varying Pruning Ratios

### v4

@@ Appendix C Details of Evaluations
Ensuring a fair and comprehensive comparison, we employed the same version of the LM Evaluation Harness as used in the SliceGPT experiments, obtaining evaluation scores under identical experimental configurations. These scores closely match those reported in the SliceGPT paper, as detailed in Table 3. For consistency, we present our reproduced results in the main experiments.
@@
For evaluating the performance of pruned models on downstream tasks, we utilized five multiple-choice QA datasets: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c. Additionally, to assess text generation quality, we calculated perplexity using the test set of the Wikitext2 dataset. For the downstream task evaluations, we adhered to the default evaluation parameters and zero-shot settings, with a batch size set to 1. For perplexity calculations, the maximum text length was set to 2048, maintaining a batch size of 1 as well.
Appendix D Perplexity and JS Divergence in Block Evaluation
Recent work such as FINERCUT (Zhang et al., 2024a) proposes a fine-grained pruning algorithm that evaluates block importance using the JS divergence between the output distributions of the original and pruned models. While this metric captures distributional shifts, it overlooks the semantic fluency and coherence of generated text—key aspects for maintaining the practical utility of LLMs. In contrast, we adopt perplexity (PPL) as a global importance metric derived from sequence-level log-likelihood, which more directly reflects how pruning impacts output quality and fluency in models.
To validate this perspective, we conducted experiments using both metrics across various model scales and pruning ratios. The results, summarized in Table 4, indicate that PPL consistently outperforms JS divergence under different configurations. These findings demonstrate that PPL better reflects the fluency and quality of the pruned model’s outputs, reinforcing its suitability as a block importance metric for LLM pruning.
Model Metric Ratio(%) Avg.Score
Llama2-7B PPL 16.98/24.99/30.00/37.02 64.33/57.65/55.01/49.36
JS 14.99/26.00/29.01/37.02 63.76/57.08/53.71/49.04
Llama2-13B PPL 14.54/25.12/31.62/34.88 69.38/64.52/61.52/60.13
JS 14.54/26.75/32.45/34.88 69.28/64.36/61.00/59.61
Table 4: Comparison of PPL and JS divergence across different pruning ratios and model scales.
Appendix E Ablation at Higher Sparsity
BlockPruner is motivated by the goal of preserving model performance more effectively through fine-grained block pruning. Evaluating how block pruning performs at different levels of granularity, particularly under higher sparsity, is crucial for supporting our motivations and claims. In light of this, we conducted ablation experiments with higher sparsity ratios on Llama2-7B and Llama2-13B models. The results, shown in Table 5, confirm that our approach remains effective, further validating the motivations behind BlockPruner.
@@ Appendix E Ablation at Higher Sparsity
Appendix F Inference Speed after Pruning
In this section, we evaluate the inference speed by measuring the time required to generate 128 tokens using models obtained from different pruning methods, all employing KV caches for efficient decoding. Each configuration is repeated 20 times to ensure statistically robust results, and we report the average inference time across runs.
As shown in Table 6, our method consistently achieves the greatest speedup at comparable pruning ratios. This improvement stems from the fact that our approach prunes a greater proportion of MHA blocks at the same overall pruning ratio compared to other methods, leading to a substantial reduction in KV cache usage, which directly accelerates the inference process of LLMs.
@@ Appendix F Inference Speed after Pruning
Appendix G Time Costs of Pruning Methods
@@
Appendix H Post-training after Pruning
We sampled 8,000 instances from the Alpaca dataset and conducted post-training on the pruned Llama2-7B and Llama2-13B models obtained via BlockPruner using LoRA. All linear layers, excluding the embedding layer and the language model head, were trained. The LoRA rank and LoRA $\alpha$ parameters were set to 32 and 10, respectively, with a learning rate of 2e-4 and a batch size of 1. Additionally, we configured the gradient accumulation steps to 4 and employed a linear learning rate scheduler. We controlled the pruning ratios within the range of 24% to 33%. The results are shown in Figure 8. It can be seen that after training, our models showed further improvement at different pruning ratios. The Llama2-7B and Llama2-13B models recovered to 89% and 92% of the performance of the unpruned models, respectively, when pruned by approximately 1/4.
@@ Appendix H Post-training after Pruning
Appendix I Sensitivity to Sample Size
@@
Appendix J Retention of Reasoning Ability After Pruning
To further assess the effectiveness of BlockPruner on reasoning-intensive scenarios, we evaluate its zero-shot performance on the AQuA-RAT dataset (Ling et al., 2017), a benchmark designed for algebraic word problem solving with rationales. We compare BlockPruner against several strong pruning baselines using the LLaMA2-7B and LLaMA2-13B models.
As summarized in Table 8, BlockPruner consistently maintains competitive performance, achieving accuracy close to the unpruned models and outperforming alternative pruning strategies such as ShortGPT, Relative Magnitude (RM), and LaCo. These results demonstrate that our method preserves the reasoning capabilities of LLMs even under significant structural compression.
Method LLaMA2-7B LLaMA2-13B
Original 31.89 30.31
BlockPruner 29.92 29.53
ShortGPT 29.13 27.95
RM 29.92 26.77
LaCo 24.80 27.95
Table 8: Zero-shot accuracy (%) on AQuA-RAT for models pruned with different methods. The pruning ratios for all methods are consistent with those used in the main experiments.
Appendix K Generalization to New Model Series
To evaluate the generalization capability of BlockPruner to newly released model series, we conduct additional experiments on two recent architectures: Mistral-7B-v0.3 (Jiang et al., 2023) and LLaMA3-8B (Grattafiori et al., 2024). We compare our method against ShortGPT and Relative Magnitude (RM) across five pruning ratios, reporting average downstream task scores on each.
As shown in Table 9, BlockPruner consistently achieves superior performance on both model families across all sparsity levels. The performance gap becomes more pronounced as the pruning ratio increases, suggesting that our fine-grained pruning approach is particularly robust under high-compression regimes. These results confirm the effectiveness and adaptability of BlockPruner beyond the models evaluated in the main paper.
Model Metric Ratio(%) Avg.Score
Mistral-7B-v0.3 BlockPruner 25.12/27.55/29.98/36.00 58.02/57.02/56.10/49.18
ShortGPT 24.07/27.08/30.09/36.11 57.40/53.26/42.79/41.41
RM 24.07/27.08/30.09/36.11 37.75/35.89/36.49/35.52
LLaMA3-8B BlockPruner 24.86/27.58/30.30/35.73 52.36/50.08/46.38/45.41
ShortGPT 24.45/27.16/29.88/35.31 41.97/44.51/43.71/42.85
RM 24.45/27.16/29.88/35.31 39.34/38.90/37.48/36.80
Table 9: Downstream task performance on recently released model families. BlockPruner consistently maintains stronger performance across pruning ratios.
Appendix L Varying Pruning Ratios
@@
Appendix M Evaluation on Additional Datasets
We extended the primary experiment by incorporating four additional well-established evaluation datasets: SWAG (Zellers et al., 2018), TruthfulQA (Lin et al., 2022), OpenBookQA (Mihaylov et al., 2018), and RACE (Lai et al., 2017). As illustrated in Table 10, our proposed method consistently surpasses previous pruning baselines across this broader range of datasets, further demonstrating its effectiveness and generalization capability.
Model Method Ratio (%) PIQA WinoGrande HellaSwag ARC-e ARC-c TruthfulQA RACE SWAG ObenBookQA Avg. Score
Llama2-7B Dense 0 79.05 69.06 75.99 74.54 46.16 25.34 39.43 76.65 44.00 58.91
SliceGPT 21.45 72.42 59.91 56.04 63.64 37.12 25.34 37.22 61.57 33.20 49.61
LaCo 21.02 68.34 60.46 54.08 55.39 35.84 28.40 29.57 61.75 39.80 48.18
RM 21.02 54.46 49.25 29.22 34.43 22.53 26.07 22.58 38.60 27.60 33.86
ShortGPT 21.02 70.24 65.90 62.63 56.06 36.09 26.81 34.07 64.84 37.20 50.43
BlockPruner 21.99 74.21 62.43 65.87 61.07 37.29 22.03 34.83 69.81 37.20 51.64
Llama2-13B Dense 0 80.52 72.14 79.36 77.36 49.23 26.07 40.77 78.04 45.40 60.99
SliceGPT 21.52 74.32 65.59 60.71 68.52 42.41 24.72 37.42 65.61 39.80 53.23
LaCo 24.37 72.42 59.27 60.44 54.34 34.56 23.62 31.87 67.93 41.00 49.49
RM 24.37 73.72 66.61 66.80 66.12 41.98 20.81 38.28 68.08 38.40 53.42
ShortGPT 24.37 72.74 70.80 67.80 60.35 41.30 24.60 37.80 68.67 41.00 53.90
BlockPruner 25.12 76.93 66.30 72.20 65.82 41.38 24.97 38.85 72.94 40.60 55.55
Baichuan2-7B Dense 0 77.48 68.27 72.18 72.98 42.75 23.01 38.76 75.26 40.00 56.74
LaCo 21.57 68.28 58.56 51.50 52.90 28.50 21.42 31.10 62.37 33.60 45.36
RM 21.57 59.96 52.33 30.87 38.17 23.63 25.09 22.01 47.38 27.40 36.32
ShortGPT 21.57 63.71 62.67 50.01 47.31 30.72 24.60 30.62 55.81 31.20 44.07
BlockPruner 22.45 69.75 61.48 58.09 58.08 33.02 20.81 33.21 64.95 32.20 47.95
Baichuan2-13B Dense 0 78.84 70.40 75.23 74.07 47.70 26.93 41.15 76.87 43.60 59.42
LaCo 22.68 70.89 58.01 54.00 57.11 32.94 20.69 29.38 67.79 33.80 47.18
RM 22.68 68.99 67.88 63.78 57.49 37.54 25.46 36.84 64.50 33.80 50.70
ShortGPT 22.68 69.31 68.27 61.71 56.52 36.69 26.93 36.27 64.14 34.40 50.47
BlockPruner 24.19 71.44 64.01 64.20 59.81 37.88 23.50 36.75 67.43 35.40 51.16
Qwen1.5-7B Dense 0 79.22 66.46 76.92 62.16 42.66 34.76 42.11 76.22 41.60 58.01
LaCo 20.97 70.40 58.64 56.35 46.89 32.85 25.34 32.92 63.43 37.20 47.11
RM 20.97 67.36 49.88 42.00 54.17 28.58 21.66 23.54 58.32 35.40 42.32
ShortGPT 20.97 69.53 62.12 58.87 43.60 32.17 32.07 31.00 57.72 33.40 46.72
BlockPruner 21.83 71.71 55.56 59.31 53.70 33.28 25.70 34.74 61.32 33.80 47.68
Qwen1.5-14B Dense 0 79.87 70.56 79.41 68.48 47.01 35.86 41.05 76.72 43.60 60.28
LaCo 22.25 71.55 58.33 60.16 53.70 34.04 22.28 33.78 65.79 35.00 48.29
RM 22.25 67.08 53.28 42.08 50.72 29.01 26.44 27.08 58.64 32.40 42.97
ShortGPT 22.25 58.60 55.96 36.16 38.09 34.81 27.05 26.99 39.89 31.40 38.77
BlockPruner 23.72 75.24 61.48 66.92 59.51 39.08 30.60 33.78 67.39 38.20 52.47
Table 10: Zero-shot downstream task performance of various models using varied pruning methods. “Dense” denotes the original, unpruned models. All evaluations are conducted with the same configuration to ensure comparability.